OpenStudy (anonymous):

A teacher collects a total of 17 phones and iPods before a group of students head on a bushwalk. From a second group of students, 40 phones and iPods are collected. The second group had tqice the number of phones and 3 times the number of iPods than the first group. How many phones and iPods did the first group have?

OpenStudy (ajprincess):

Let us assume that number of ipods the first group had be x nd number of phones be y. then x+y=17 3x+2y=40 Can u solve these simultaneous equations? @oly.surf

OpenStudy (texaschic101):

it would be easier to solve by substitution :)

OpenStudy (linyu):

lets the first group be ax+by=17 second group be Ax+By=40 The sum of the whole groups (A+a)x+(B+b)y=57 We have a relation A/a=2 as to A=2a, B/b=3 as to B=3b Therefore the total sum can be written as (2a+a)x+(3b+b)y=57 3ax+4by=57. a and b are just expressions for the relationship of first and second group so we delete them and write the first group and the sum equation x+y=17 3x+4y=57 can you solve it?

OpenStudy (mayankdevnani):

by moving towards the @ajprincess method.. x+y=17 3x+2y=40 First solve. x+y=17 x=17-y and then put x=17-y in 3x+2y=40 3(17-y)+2y=40 can you solve it???? @oly.surf

OpenStudy (mayankdevnani):